Find the position vector of the mid point of the vector joining the points P (2, 3, 4) and Q (4, 1, – 2).

Grade:

2 Answers

Latika Leekha
9 years ago
The given points are P (2, 3, 4) and Q(4, 1, -2).
Now, Let OS be the position vector of the mid point joining these points.
(The vectors have been denoted by bold letters)
P can be written as 2i + 3j + 4k
Q can be written as 4i + j – 2k
Now, the coordinates of the reqiured position vector are gievn by
[(2i + 3j + 4k) + (4i + j – 2k)]/2
= [(2i + 4i) + (3j + j) + (4k– 2k)] / 2
= (6i + 4j + 2k) / 2
= 3i + 2j + k
Thanks & Regards
Latika Leekha
askIITians Faculty
Rahul
6 years ago
Vector P = 2i+3j+4kVector Q=4i+j-2k Required Position vector = (vector P +vector Q)/2=> 6i+4j+2k/2 = 3i+2j+k